A random sample of 30 hotels and motels on a given Sunday in Monterey found the mean to be $100 with population standard deviation $30. Find a 90% confidence interval for the true mean
Answer #1

Solution Given that Ž = $100 6 = $ 30 n = 30 At got confidence level z is d:1-904 =1-0.9 = 0 a ool 0.05 2 - 2 - 2 x 12 = 70.o